solve this equation for x
soooooooo confused
(1+2i)x+3-2i = 4-5i+3ix

Answer by sudhanshu_kmr(1152)


(1+2i)x+3-2i = 4-5i+3ix
=> x + 2xi + 3 -2i = 4 -5i + 3xi
=> x + 2xi -3xi = 4 -5i -3 +2i
=> x -xi = 1 -3i
=> x(1-i) = 1-3i
=> x = (1-3i)/(1-i)
you can also simplify - multiplying nominator and denominator by (1+i)
it will give
x = 2 - i